# MambaTab: A Simple Yet Effective Approach for Handling Tabular Data

### v1

@@ 4.2 Vanilla Supervised Learning Performance
4.3 Feature Incremental Learning Performance
For this setting, we divide the feature set $F$ of each dataset into three non-overlapping subsets $s_{1},s_{2},s_{3}$. $set_{1}$ contains $s_{1}$ features, $set_{2}$ contains $s_{1},s_{2}$ features, and $set_{3}$ contains all features in $s_{1},s_{2},s_{3}$. While other baselines can only learn from either $set_{1}$ by dropping all incrementally added features (with respect to $s_{1}$) or $set_{3}$ by dropping old data, TransTab Wang and Sun (2022) and MambaTab can incrementally learn from $set_{1}$ to $set_{2}$ to $set_{3}$. In our method, we simply change the input feature cardinality $n(set_{i})$ between settings, with the architecture fixed. Our method works because Mamba has strong content and context selectivity for extrapolation and we keep the representation space dimension fixed, that is, independent of feature set cardinality $n(F)$. Thus, this demonstrates the adaptability and simplicity of our method for incremental environments. Even with default hyperparameters, MambaTab-D outperforms all baselines as shown in Table 3. Here, we report the results averaged over 10 runs with different random seeds. Since it already achieves strong performance, we do not tune the hyperparameters further, although doing so could potentially improve performance.
Table 3: Test AUROC results on 8 public datasets for feature incremental learning. Results reported here are averaged for 10 runs with random splits for our method MambaTab-D. The best results are shown in bold.
@@ 4.3 Feature Incremental Learning Performance
4.4 Learnable Parameter Comparison
Our method not only achieves superior performance compared to existing state-of-the-art methods, it is also memory and space efficient. We demonstrate our method’s superiority in terms of learnable parameter size while comparing against transformer-based approaches in Table 4. It is seen that our method (both Mambda-D and -T) achieves comparable or better performance than TransTab typically with $<1\%$ of its learnable parameters. To evaluate learnable parameter size, we use the default settings specified in FT-Trans, TransTab, and TabTrans  [^2] https://github.com/lucidrains/tab-transformer-pytorch  [^3] https://github.com/RyanWangZf/transtab. We also notice that, despite varying features, TransTab’s model size remains unchanged. The most important tunable hyperparameters for MambaTab include the block expansion factor (the local kernel size), the state expansion factor ($N$), and the embedded representation space dimension. We perform sensitivity analysis on them in Section 5 and also fine-tune them for each dataset. In addition, we conduct ablation study for the normalization layer of our model.
Table 4: Comparison of total learnable parameters between our method MambaTab and transformer-based methods. (M = million, K = thousand)
@@ 4.4 Learnable Parameter Comparison
4.5 Hyperparameter Tuning
As mentioned above, we tune important hyperparameters and use the validation loss for tuning. Therefore, the test set is never used for tuning. The hyperparameters achieving the best validation loss are used for testing. We have reported averaged test results over 10 runs with different random seeds with the tuned MambaTab (Table 2). Learnable parameter sizes of MambaTab-T are reported in Table 4. Interestingly, MambaTab-T sometimes consumes fewer parameters than even MambaTab-D, e.g., on DS, BL, IO, and IC. We demonstrate key components of our tuned model MambaTab-T in Table 5, where the tuned values for these components are shown, with other training-related hyperparameters, such as batch size and learning rate, at default values; see Implementation Details in Section 4.
Table 5: Hyperparameters of our tuned model, MambaTab-T. The performance of MambaTab-T for vanilla supervised learning has been shown in Table 2.
@@ 4.5 Hyperparameter Tuning
5 Hyperparameter Sensitivity Analysis and Ablation Study
@@
5.1 Block Expansion Factor
We experiment with block expansion factor (kernel size) $\{1,2,...,10\}$, keeping the other hyperparameters at default values as in MambaTab-D. As seen in Figure 3, MambaTab’s performance changes only slightly with different block expansion factors, with no clear or monotonic trends. Thus we set the default to 2, inspired by Gu and Dao (2023), though tuning this parameter further could improve performance on some datasets.
Figure 3: Sensitivity of MambaTab AUROC to block expansion factor on two randomly chosen datasets CB and CG.
5.2 State Expansion Factor
We demonstrate the effect of the state expansion factor ($N$) using values in $\{4,8,16,32,64,128\}$. As seen in Figure 4, MambaTab performance in AUROC improves with increasing state expansion factor for both datasets CG and CB. Thus tuning this hyperparameter could further improve performance. However, a larger state expansion factor consumes more memory. To balance performance versus memory consumption, we select 32 as the default value.
Figure 4: Sensitivity of MambaTab AUROC to state expansion factor on two randomly chosen datasets CB and CG.
5.3 Size of Embedded Representations
@@
Figure 5: Sensitivity of MambaTab performance in AUROC to size of embedded representations on two datasets CB and CG.
5.4 Ablation of Layer Normalization
We demonstrate the effect of layer normalization, which is applied to the embedded representations, in our model architecture shown in Figure 1. We contrast the performance by keeping or dropping this layer in vanilla supervised learning experiments on CG and CB datasets. The results in AUROC metric are shown in Table 6.
Table 6: Ablation analysis of layer normalization. Experiments are conducted with and without layer normalization under supervised learning using the model architecture shown in Figure 1. Results demonstrated are on test set AUROC.
@@ 5.4 Ablation of Layer Normalization
We can see the effectiveness of the normalization layer. Without layer normalization, the embeddings would directly pass through the ReLU activation, as shown in the overall scheme (Figure 1). On both CG and CB datasets, MambaTab’s performance improves with layer normalization versus without. This ablation thus justifies the incorporation of layer normalization in our model.
5.5 Effect of Batch Size
In addition to the above model-related hyperparameters, we also perform sensitivity analysis with experiments on batch size. Due to its small model size, MambaTab can easily handle a large number of samples per batch. We demonstrate AUROC results in Figure 6 using sizes of $\{60,80,100,120,140\}$. We see small variations in performance for both CG and CB datasets across minibatch sizes. This demonstrates our method’s generalization capability regarding batch size. Considering this insensitivity, we set 100 as the default batch size.
Figure 6: Sensitivity of MambaTab performance in AUROC to batch size. Other parameters are kept at default values while the batch size is varied from 60 to 140 in steps of 20.
5.6 Scaling Mamba
@@
Figure 7: AUROC results and learnable parameter sizes versus number of stacked residual Mamba blocks ($\mathcal{M}$) on CB and CG datasets. Other parameters are kept at default values of MambaTab-D.
Although we have achieved comparable or superior performance to current state-of-the-art methods with a default ${\mathcal{M}}=1$ under regular supervised learning (see Table 2), we also study the effect of scaling Mamba blocks via residual connections following  He et al. (2016). We stack Mamba blocks as in Figure 1, concatenating $\mathcal{M}=2$ up to 100 blocks, as shown in Equation 5:
@@
Here, $h^{(i)}$ is the hidden state from the $i$-th Mamba block, that is, $Mamba_{i}$, taking the prior block’s hidden state $h^{(i-1)}$ as input. As seen in Figure 7, with increasing Mamba blocks, MambaTab retains comparable performance while the learnable parameters increase linearly on both CG and CB datasets. This demonstrates Mamba block’s information retention capacity. We observe that few Mamba blocks suffice for strong performance. Hence, we choose to use $\mathcal{M}=1$ by default in MambaTab.
6 Future Scope
@@
7 Conclusion
This paper presents Mambatab, a simple yet effective method for handling tabular data. It uses Mamba, a state-space-model variant, as a building block to classify tabular data. MambaTab can effectively learn and predict in both vanilla supervised learning and feature incremental learning settings. Moreover, it requires no manual data preprocessing. MambaTab demonstrates superior performance over current state-of-the-art deep learning and traditional machine learning-based baselines under both supervised and incremental learning on 8 public datasets. Remarkably, MambaTab occupies only a small fraction of memory in learnable parameter size compared to Transformer-based baselines for tabular data. Extensive results demonstrate MambaTab’s efficacy, efficiency, and generalizability for diverse tabular learning applications.
@@ Acknowledgments
We thank the creators of the public datasets used, for making these valuable resources openly available for research purposes. We also thank the authors of the baseline models and Mamba Gu and Dao (2023) for releasing their valuable code publicly. This research is supported in part by the NSF under Grant IIS 2327113 and the NIH under Grants R21AG070909, P30AG072946, and R01HD101508-01.
Supplementary Materials
Dataset details with accessible links
Here in this section, we provide the dataset accessible links with their abbreviations.
Table 1: Publicly available datasets abbreviations and external accessible link
Instructions on how to run the code
The following instructions are helpful to run our codes. The instructions are written in a sequential manner.
Installation
Please install the following required libraries:
  • pip install torch==2.1.1 torchvision==0.16.1
  • pip install causal-conv1d==1.1.1
  • pip install mamba-ssm
File instruction and brief definition
  • $\bf config.py$: contains training related configurations setting.
  • $\bf MambaTab.py$: contains our method-related code of MambaTab.
  • $\bf supervised\_mambatab.py$: contains code related to vanilla supervised learning setting.
  • $\bf feature\_incremental.py$: contains code related to incremental feature setting.
  • $\bf train\_val.py$: contains training and validation related code for training and validating the model over the epochs.
  • $\bf utility.py$: contains code for data reading and data pre-processing.
Data download and processing
Please download data using the accessible links mentioned in the supplementary materials table  1.
Cautions:
  • Dataset must be in .$csv$ format.
  • Header row must be in the first row in the .$csv$
  • Target column must be in the last column in the .$csv$
  • Rename the $X.csv$ to $data\_processed.csv$ and place it into $datasets/X$ folder. Here ’X’ can be ’dress’, ’cylinder’, etc.
Configurations
Please utilize our $\bf config.py$ file for the necessary configurations to run the code. Moreover, more configurations can also be modified in our $\bf MambaTab.py$ file for model-related configurations.
Running specific files
  • To run vanilla supervised learning, please execute $\bf supervised\_mambatab.py$. Please make sure to change the necessary configurations in $\bf config.py$ before running this file. We have mentioned the necessary comments for an easier understanding of the flow of the code.
  • To run the feature incremental learning setting, please execute $\bf feature\_incremental.py$. Here, as mentioned in our paper, we divide the features into 3 non-overlapping sets and perform the training incrementally.
  • As mentioned above, our code is flexible in changing parameters and getting results with tuned settings.

### v2

@@ IV-B Vanilla Supervised Learning Performance
IV-C Feature Incremental Learning Performance
For this setting, we divide the feature set $F$ of each dataset into three non-overlapping subsets $s_{1},s_{2},s_{3}$. $set_{1}$ contains $s_{1}$ features, $set_{2}$ contains $s_{1},s_{2}$ features, and $set_{3}$ contains all features in $s_{1},s_{2},s_{3}$. While other baselines can only learn from either $set_{1}$ by dropping all incrementally added features (with respect to $s_{1}$) or $set_{3}$ by dropping old data, TransTab [2] and MambaTab can incrementally learn from $set_{1}$ to $set_{2}$ to $set_{3}$. In our method, we simply change the input feature cardinality $n(set_{i})$ between settings, with the remaining architecture fixed. Our method works because Mamba has strong content and context selectivity for extrapolation and we keep the representation space dimension fixed, that is, independent of feature set cardinality $n(F)$. Thus, this demonstrates the adaptability and simplicity of our method for incremental environments. Even with default hyperparameters, MambaTab-D outperforms all baselines as shown in Table III. We could not directly find the performance of UniTabE for the feature incremental learning setting. Here, we report the results averaged over 10 runs with different random seeds. Since it already achieves strong performance, we do not tune the hyperparameters further, although doing so could potentially improve performance.
TABLE III: Test AUROC for feature incremental learning. The best results are shown in bold.
@@ IV-C Feature Incremental Learning Performance
IV-D Self-Supervised Leraning
For self-supervised learning (SSL), we randomly corrupted 50% features in each iteration with zeros and tried to reconstruct the original sequence with our model. To achieve this, we utilized $L_{2}$ loss for reconstruction and changed the output projection dimension to be matched with the number of features passed as input to our model. With this strategy, our model learns more impactful underlying meanings of the data. Here again, we only utilized the train-val set for self-supervised learning keeping the test set completely unseen to the model. Moreover, our SSL technique is completely unaware of the true class label of the feature and only utilizes raw features, demonstrating its robustness towards leveraging large-scale unlabeled data. Our performance compared against TransTab’s two pre-training strategies of Vertical-Partition Contrastive Learning (VPCL) via supervised and self-supervised techniques are shown in Table IV, which demonstrates our method’s superiority over TransTab while performing SSL fashioned learning. Moreover, our method also demonstrates superior performance over UniTabE-FT in many cases.
TABLE IV: Test AUROC results under different pre-training schemas.
Methods Schema Datasets
CG CA DS AD CB BL IO IC
TransTab Self-VPCL 0.777 0.837 0.626 0.907 0.819 0.843 0.823 0.919
VPCL 0.776 0.858 0.637 0.907 0.862 0.844 0.819 0.919
UniTabE FT 0.790 0.940 0.660 0.910 0.880 0.840 0.760
MambaTab SSL 0.804 0.967 0.649 0.909 0.880 0.857 0.786 0.909
IV-E Learnable Parameter Comparison
Our method not only achieves superior performance compared to existing state-of-the-art methods, it is also memory and space efficient. We demonstrate our method’s superiority in terms of learnable parameter size while comparing against transformer-based approaches in Table V. It is seen that our method (both MambaTab-D/T) achieves comparable or better performance than TransTab typically with $<1\%$ of its learnable parameters. To evaluate learnable parameter size, we use the default settings specified in FT-Trans, TransTab, and TabTrans  [^2] https://github.com/lucidrains/tab-transformer-pytorch. We also notice that, despite varying features, TransTab’s model size remains unchanged. The most important tunable hyperparameters for MambaTab include the controllable block expansion factor, the state expansion factor ($N$), and the embedded representation space dimension. We perform sensitivity analysis on them in Section V and also fine-tune them for each dataset. In addition, we conduct an ablation study for the normalization layer of our model.
TABLE V: Total learnable parameters (M = million, K = thousand).
@@ IV-E Learnable Parameter Comparison
IV-F Hyperparameter Tuning
We tune critical hyperparameters using validation loss, ensuring the test set remains untapped until final testing with the best-performing settings. We have reported averaged test results over 10 runs with different random seeds with the tuned MambaTab (Table II). Learnable parameter sizes of MambaTab-T are reported in Table V. Interestingly, MambaTab-T sometimes consumes fewer parameters than even MambaTab-D, e.g., on DS, BL, IO, and IC. We demonstrate key components of our tuned model MambaTab-T in Table VI, where the tuned values for these components are shown, with other training-related hyperparameters, such as training epochs and learning rate, at default values; see Implementation Details in Section IV.
TABLE VI: Hyperparameters of our tuned model, MambaTab-T.
@@ IV-F Hyperparameter Tuning
V Hyperparameter Sensitivity and Ablation Study
@@
V-A Block Expansion Factor
We experiment with block expansion factor (kernel size) $\{1,2,...,10\}$, keeping the other hyperparameters at default values as in MambaTab-D. As seen in Figure 3, MambaTab’s performance changes only slightly with different block expansion factors, with no clear or monotonic trends. Thus we set the default to 2, inspired by [6], though tuning this parameter further could improve performance on some datasets.
V-B State Expansion Factor
We show the impact of the state expansion factor ($N$) using values in ${4,8,16,32,64,128}$, where MambaTab’s AUROC improves with increasing $N$ for datasets CG and CB (Figure 3). Although a higher $N$ enhances performance, it also uses more memory. Balancing performance and memory use, we set 32 as the default $N$.
(a) Block Expansion Factor
(b) State Expansion Factor
Figure 3: Ablation on block and state expansion factors.
V-C Size of Embedded Representations
@@
Figure 4: Ablation on embedded representation space.
V-D Ablation of Layer Normalization
We demonstrate the effect of layer normalization, which is applied to the embedded representations, in our model architecture shown in Figure 1. We contrast the performance by keeping or dropping this layer in vanilla supervised learning experiments on CG and CB datasets. The results in AUROC metric are shown in Table VII.
TABLE VII: Ablation analysis of layer normalization via AUROC
@@ V-D Ablation of Layer Normalization
Without layer normalization, the embeddings would directly pass through the ReLU activation, as shown in the overall scheme (Figure 1). On both CG and CB datasets, MambaTab’s performance improves with layer normalization versus without.
V-E Scaling Mamba
@@
Figure 5: Analysis on the of stacked residual Mamba blocks ($\mathcal{M}$).
Although we have achieved comparable or superior performance to current state-of-the-art methods with a default ${\mathcal{M}}=1$ under regular supervised learning (see Table II), we also study the effect of scaling Mamba blocks via residual connections following  [22]. We stack Mamba blocks as in Figure 1, concatenating $\mathcal{M}=2$ up to 100 blocks, as shown in Equation 3:
@@
Here, $h^{(i)}$ is the hidden state from the $i$-th Mamba block, that is, $Mamba_{i}$, taking the prior block’s hidden state $h^{(i-1)}$ as input. As seen in Figure 5, with increasing Mamba blocks, MambaTab retains comparable performance while the learnable parameters increase linearly on both CG and CB datasets. This demonstrates the Mamba block’s information retention capacity. We observe that few Mamba blocks suffice for strong performance. Hence, we choose to use $\mathcal{M}=1$ by default.
VI Future Scope
@@
VII Conclusion
This paper presents MambaTab, a plug-and-play method for learning tabular data. It uses Mamba, a state-space-model variant, as a building block to classify tabular data. MambaTab can effectively learn and predict in vanilla supervised learning, feature incremental learning settings, and adaptable to self-supervised learning. MambaTab demonstrates superior performance over current state-of-the-art deep learning and traditional machine learning-based baselines under supervised, feature incremental learning, and self-supervised learning on 8 public benchmark datasets. Remarkably, MambaTab occupies only a small fraction of memory in learnable parameter size compared to Transformer-based baselines for tabular data. Extensive results demonstrate MambaTab’s efficacy, efficiency, and generalizability across diverse datasets for various tabular learning applications.
@@ Acknowledgments
We thank the creators of the public datasets and the authors of the baseline models and Mamba [6] for making these resources available for research. This research is supported in part by the NSF under Grant IIS 2327113 and the NIH under Grants R21AG070909, P30AG072946, and R01HD101508-01.